# Refining Concentration for Gaussian Quadratic Chaos

### v2

@@ Abstract
We visit and slightly modify the proof of Hanson-Wright (HW) inequality for concentration of Gaussian quadratic chaos where we are able to tighten the bound by increasing the absolute constant in its formulation from its largest currently known value of 0.125 to at least 0.145 in the general symmetric case. We also present a sharper version of an inequality due to Laurent and Massart (LM inequality) through which we are able to increase the absolute constant in HW inequality from its largest currently available value of approximately $0.134$ due to LM inequality itself to at least $0.152$ in the positive-semidefinite case. A new sequence of concentration bounds indexed by $m=1,2,3,\cdots$ is developed for Gaussian quadratic chaos that involve Schatten norms of the underlying matrix. The case $m=1$ recovers HW inequality. These bounds undergo a phase transition in the sense that if the tail parameter is smaller than a critical threshold, then the case $m=1$ (HW inequality) is the tightest and if the tail parameter is larger than this threshold, then the bounds eventually become tighter than any bound of finite index as $m$ grows to infinity. This leads to what is referred to as the $m_{\infty}$-bound. An estimate (upper bound) is provided for the aforementioned threshold. We also develop a concentration bound that is tighter than HW inequality for both sufficiently small and large values of the tail parameter. Finally, we explore concentration bounds when the underlying matrix is positive-semidefinite and only the dimension $n$ and its operator norm (largest eigenvalue) are known. Four candidates are examined, namely, the $m_{\infty}$-bound, relaxed versions of HW and LM bounds and a bound that we refer to as the $\chi^{2}$-bound. It is proved that the sharpest among these is always either the $m_{\infty}$-bound or the $\chi^{2}$-bound. The case of even dimension is given special attention. If $n=2,4,6$, the $\chi^{2}$-bound is always tighter than the $m_{\infty}$-bound. If $n$ is an even integer greater than or equal to $8$, it is established that the $m_{\infty}$-bound is sharper than the $\chi^{2}$-bound if and only if the ratio of the tail parameter over the largest eigenvalue lies inside a nonempty finite open interval which expands indefinitely as $n$ grows.
Index Terms:
HW inequality, LM inequality, $m_{\infty}$-bound, $\chi^{2}$-bound, Descartes’ Rule of Signs.
@@ I-A Summary of prior art
This paper concerns the Gaussian version of Hanson-Wright (HW) inequality [1], a concentration of measure result that presents an upper bound on the tail probability for a quadratic form in a random vector of independent standard Gaussian random variables also known as order 2 chaos. [^1] In its most general form, HW inequality concerns quadratic forms in sub-Gaussian random vectors. For an integer $n\geq 1$ called the dimension throughout the paper, let $\underline{\boldsymbol{x}}$ be such a random vector of length $n$ and $A$ be an arbitrary $n\times n$ matrix. Denote the difference between $\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}$ and its expected value $\mathbb{E}[\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}]=\mathrm{Tr}(A)$ by $\boldsymbol{\Delta}$, i.e.,
@@
HW inequality states that for every tail parameter $t>0$,
@@
where $\|A\|_{2}$ is the Hilbert-Schmidt norm of $A$, $\|A\|$ is the operator norm of $A$ and $\kappa$ is an absolute constant that does not depend on $n$, $A$ and $t$ [^2] Replacing $A$ by $-A$ in (2) results in the same upper bound as in (2) on the left-tail probability $\Pr(\boldsymbol{\Delta}<-t)$. Without loss of generality, we assume $A$ is symmetric throughout the paper. This can be easily justified by noting that $\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}=\underline{\boldsymbol{x}}^{T}\tilde{A}\underline{\boldsymbol{x}}$ where $\tilde{A}=\frac{1}{2}(A+A^{T})$ is symmetric and $\|\tilde{A}\|_{2}\leq\|A\|_{2}$ and $\|\tilde{A}\|\leq\|A\|$. These inequalities follow by applying the triangle inequality for norms and noting that $\|A\|_{2}=\|A^{T}\|_{2}$ and $\|A\|=\|A^{T}\|$. In the original paper [1], $\|A\|$ is the operator norm of the matrix whose entries are absolute values of entries of $A$. Reference [2] shows that the inequality indeed holds as stated in (2) with $\|A\|$ being the operator norm of $A$ itself. An explicit value for $\kappa$ is missing or at least hard to identify in [1, 2]. Recently, Textbook[3] gives the explicit value $\kappa=\frac{1}{8}=0.125$ at the end of the proof for Theorem B.8 on page 304. This value for $\kappa$ can also be inferred form inequality (3.28) on page 120 in [4] given by
@@
It is stated in [4] that (3) holds for arbitrary symmetric matrix $A$. However, there is a mistake in the proof. Inequality $\frac{1}{2}(-\ln(1-2x)-2x)\leq\frac{x^{2}}{1-2x}$ in the middle of page 120 does not hold for $x<0$. Nonetheless, inequality (3) certainly holds when $A$ is positive-semidefinite. Loosening the bound by writing $\|A\|_{2}^{2}+\|A\|t\leq 2\max\{\|A\|_{2}^{2},\|A\|t\}$, we arrive at HW inequality with $\kappa=0.125$. For a positive-semidefinite matrix $A$, a key result in the literature on concentration of Gaussian quadratic chaos is inequality (4.1) on page 1325 in [5] due to Laurent and Massart given by
@@
for all tail parameters $t>0$. We will refer to (4) as Laurent-Massart (LM) inequality. One can show that LM inequality is equivalent to [^3] See Section V.
@@
We will show in Section V that LM inequality implies HW inequality at best with $\kappa=1-\frac{\sqrt{3}}{2}\approx 0.134$.
@@ I-B Contributions and main results
The main contributions in this paper are fivefold:
@@
    We visit and slightly modify the proof given in [3] for HW inequality. In the course of proof, Reference [3] uses the bound
@@
    for $a,b>0$. For given $10<b<$ 1, we determine the smallest $a$ (tightest bound) such that (7) holds. As a result, we are able to increase $\kappa$ from $0.125$ to at least $0.145$ for arbitrary symmetric matrix $A$. More precisely, we have the following proposition.
@@
    For $10<b<$ 1 and integer $m\geq 1$ define the sequence of functions
    $\theta_{m}(b)=\sum_{i=0}^{\infty}\frac{b^{i}}{i+m+1}.$ (8)
    If $A$ is symmetric, then HW inequality holds with
@@
    where $b^{*}$ is the unique solution for $b$ in the equation $12b\theta_{1}(b)=$ 1.
@@
    In the positive-semidefinite case, we present a sharper version of LM inequality. To prove (4), Reference [5] relies on the bound
[…]
@@
[…]
    for $a,b>0$. We show that for every $\frac{2}{3}<a\leq 1$, this inequality holds for $b$ at least as large as $\frac{3a-2}{a(2a-1)}$. As a result, we derive an improved LM inequality as presented in the next proposition.
@@
    Choosing $a=b=1$ and $c=\infty$ in (14) recovers LM inequality in (5). In Subsection II-A, we minimize the upper bound on the right side of (14) in terms of $a$ and compare the resulting optimized inequality with LM inequality to demonstrate the improvement. We will see that the minimizing value of $a$ is the unique root inside the interval $(\frac{2}{3},1)$ for a quintic polynomial equation [^5] See Equation (42). and it depends on the matrix $A$ and the tail parameter $t$ only through the ratio $\rho=\frac{t\|A\|}{\|A\|_{2}^{2}}$. We are able to approximate this unique root by an analytic formula [^6] See Equation (37). where the approximation error is shown to be less that $0.035$ regardless of $A$ and $t$. As a consequence of Proposition 2, we offer the next result which is a stronger version of Proposition 1 in the positive-semidefinite case.
@@
    Let $A$ be positive-semidefinite. The inequality in (14) implies HW inequality at best with
@@
    Moreover, LM inequality in (4) implies HW inequality at best with $\kappa=1-\frac{\sqrt{3}}{2}\approx 0.134$.
@@
    We explore beyond HW inequality for arbitrary symmetric matrix $A$ by considering an extension of (7), i.e., an inequality of the form
@@
    for a given integer $m\geq 1$ and $a,b>0$. We show [^7] See Lemma 1 in Section III. that for every $m\geq 1$ and $10<b<$ 1, the smallest $a$ (tightest upper bound) for which (20) holds is $\theta_{m}(b)$ as defined in (8). Consequently, we obtain a sequence of concentration bounds for Gaussian quadratic chaos that are written in terms of Schatten norms of $A$ [^8] See Subsection I.C for the definition of Schatten $p$-norm.
    Proposition 4.
    Let $m\geq 1$ be an integer, $10<b<$ 1 and $\|A\|_{m+1}$ be Schatten norm of order $m+1$ for a symmetric matrix $A$. Then
@@
    Choosing $m=1$ in (21) recovers HW inequality in (2) with $\kappa=\kappa_{1}(b)$. Schatten norm $\|A\|_{m+1}$ should not be mistaken with the $L_{p}$-norm of $A$ with $p=m+1$ unless $m=1$ in which case $\|A\|_{2}$ is Hilbert-Schmidt norm of $A$. The upper bound $e^{\Lambda_{m}(t,b)}$ in (21) is to be minimized over $10<b<$ 1.101010The minimum exists. In general, the sequence of upper bounds $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ for $m=1,2,3,\cdots$ is not monotone in $m$ for given $A$ and $t$. However, we conjecture that it undergoes a phase transition as stated next.
@@
    For a given symmetric matrix $A$, there exists a critical threshold $t_{c}>0$ such that if $t<t_{c}$, then $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ is the smallest for $m=1$ and if $t>t_{c}$, then $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ is the smallest when $m$ grows to infinity. [^11] More precisely, for $t>t_{c}$, the limit $\lim_{m\to\infty}\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ exists and it is smaller than $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ for every $m\geq 1$.
    In Subsection II-B, we will study the bounds in (21) both through numerical examples and analytically where three corollaries of Proposition 4 are stated. The first one offers loosened versions of the upper bounds in (21) which are more convenient to use as they do not involve minimization over $b$ anymore. The second corollary looks into the limiting value of (21) as $m$ grows large. This leads to the simple bound
@@
    which we refer to as the $m_{\infty}$-bound or $m_{\infty}$-inequality. The right hand side in (24) depends on $A$ only through its dimension $n$ and its operator norm $\|A\|$. The third corollary presents an estimate (upper bound) for the critical threshold $t_{c}$ mentioned in Conjecture 1 in above.
@@
    The upper bounds in Proposition 4 become tighter than the upper bound in HW inequality only when $t>t_{c}$, i.e., for all sufficiently large values of $t$. We raise the question whether one can improve over HW inequality for all sufficiently small values of $t$. We answer this question in the affirmative. We are able to present what we refer to as a “twin inequality” to HW inequality which is tighter than HW inequality not only for all sufficiently small $t$, but also for all sufficiently large $t$. This bound depends on all norms $\|A\|,\|A\|_{2}$ and $\|A\|_{3}$.
    Proposition 5.
@@
    The bound in (28) is named a twin to HW inequality in the sense that $\min\{\eta_{1}(t),\eta_{2}(t)\}=\frac{t^{2}}{\|A\|_{2}^{2}}+O(t^{3})$ as $t\to 0$ and $\min\{\eta_{1}(t),\eta_{2}(t)\}=\frac{t}{\|A\|}+O(1)$ as $t\to\infty$. These statements are verified in Subsection II-C. Hence, the inequality in (28) is a look-alike to HW inequality both as $t\to 0$ and $t\to\infty$. Note that the constant $\kappa^{\prime}\approx 0.187$ is slightly larger than the constant $\kappa\approx 0.145$ in HW inequality. This is exactly why the twin bound is sharper than HW inequality for both small and large values of $t$.
@@
    And lastly, we raise and attempt to answer the following question in Subsection II-D.
    Let $A$ be an unknown $n\times n$ positive-semidefinite matrix for which we only know the dimension $n$ and the value of its operator norm $\|A\|$ denoted by $\nu$. How much is $\sup_{A\in\mathbb{S}^{n}_{+}:\|A\|=\nu}\Pr(\boldsymbol{\Delta}>t)$ in terms of $n,\nu,t$? Here, $\mathbb{S}^{n}_{+}$ is the cone of $n\times n$ positive-semidefinite matrices.
    We present four bounds on $\Pr(\boldsymbol{\Delta}>t)$ that depend on $A$ only through $n$ and $\|A\|$. One candidate is the $m_{\infty}$-bound in (24). The second and third candidates are relaxed versions of HW and LM inequalities, respectively, which do not depend on $\|A\|_{2}$ anymore. The relaxed HW inequality is
@@
    where $F_{\chi^{2}_{n}}(\cdot)$ is the cumulative distribution function (CDF) for a $\chi^{2}$ random variable with $n$ degrees of freedom. Derivation of the relaxed HW and LM inequalities and the $\chi^{2}$-bound is presented in Subsection II-D. Note that all of these bounds in (24), (29), (30) and (31) depend on $A$ and $t$ only through the dimension $n$ and the ratio $\frac{t}{\|A\|}$. Our first result is the following.
@@
    The $m_{\infty}$-bound in (24) is sharper than both the relaxed HW inequality in (29) and the relaxed LM inequality in (30) for every positive-semidefinite matrix $A$ and $t>0$.
@@
    It follows from Proposition 6 that among the four candidates we have introduced, the sharpest bound is always either the $m_{\infty}$-bound or the $\chi^{2}$-bound. Our last proposition compares these two finalists when the dimension $n$ is even.
@@
    We close this section by mentioning that a loosened version of the $\chi^{2}$-bound in (31) known as the large deviation bound is always looser than the $m_{\infty}$-bound for every $n\geq 2$. This is shown at the end of Subsection II-D

### v3

@@ Abstract
We start by slightly modifying the proof of Hanson-Wright inequality (HWI) for concentration of Gaussian quadratic chaos where we tighten the bound by increasing the absolute constant in its formulation from the largest known value of 0.125 to at least 0.145 in the symmetric case. We also present a sharper version of an inequality due to Laurent and Massart (LMI) through which we increase the absolute constant in HWI from the largest available value of approximately $0.134$ due to LMI itself to at least $0.152$ in the positive-semidefinite case. A new sequence of concentration bounds indexed by $m=1,2,3,\cdots,\infty$ is developed that involves Schatten norms of the underlying matrix. The case $m=1$ recovers HWI. These bounds undergo a phase transition in the sense that if the tail parameter is smaller than a critical threshold $\tau_{c}$, then $m=1$ is the tightest and if it is larger than $\tau_{c}$, then $m=\infty$ is the tightest. This leads to a novel and simple bound referred to as the $m_{\infty}$-bound that depends on the underlying matrix only through its operator norm. A separate concentration bound named “twin to HWI” is also developed that is tighter than HWI for both sufficiently small and large values of the tail parameter. The twin bound involves Schatten norms of orders $2$ and $3$ and the operator norm. To further investigate the $m_{\infty}$-bound, we explore concentration bounds when the underlying matrix is positive-semidefinite and only the dimension $n$ and its largest eigenvalue are known. Five candidates are examined, namely, the $m_{\infty}$-bound, relaxed versions of HWI and LMI, the $\chi^{2}$-bound and the large deviations bound. The sharpest among these is always either the $m_{\infty}$-bound or the $\chi^{2}$-bound. The case of even dimension is given special attention. If $n=2,4,6$, the $\chi^{2}$-bound is tighter than the $m_{\infty}$-bound. If $n$ is an even integer greater than or equal to $8$, the $m_{\infty}$-bound is sharper than the $\chi^{2}$-bound if and only if the ratio of the tail parameter over the largest eigenvalue lies inside a finite open interval which expands indefinitely as $n$ grows.
@@ I-A Summary of prior art
This paper concerns the Gaussian version of Hanson-Wright inequality (HWI) [1], a concentration of measure result that presents an upper bound on the tail probability for a quadratic form in a random vector of independent standard Gaussian random variables also known as order 2 chaos. [^1] In its most general form, HWI concerns quadratic forms in sub-Gaussian random vectors. For an integer $n\geq 1$ called the dimension throughout the paper, let $\underline{\boldsymbol{x}}$ be such a random vector of length $n$ and $A$ be an arbitrary $n\times n$ matrix. Denote the difference between $\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}$ and its expected value $\mathbb{E}[\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}]=\mathrm{Tr}(A)$ by $\boldsymbol{\Delta}$, i.e.,
@@
HWI states that for every tail parameter $t>0$,
@@
where $\|A\|_{2}$ is the Hilbert-Schmidt norm of $A$, $\|A\|$ is the operator norm of $A$ and $\kappa$ is an absolute constant that does not depend on $n$, AA and tt. [^2] Replacing AA by A-A in (2) results in the same upper bound as in (2) on the left-tail probability $\Pr(\boldsymbol{\Delta}<-t)$. Without loss of generality, we assume $A$ is symmetric throughout the paper. This can be easily justified by noting that $\underline{\boldsymbol{x}}^{T}A\underline{\boldsymbol{x}}=\underline{\boldsymbol{x}}^{T}\tilde{A}\underline{\boldsymbol{x}}$ where $\tilde{A}=\frac{1}{2}(A+A^{T})$ is symmetric and $\|\tilde{A}\|_{2}\leq\|A\|_{2}$ and $\|\tilde{A}\|\leq\|A\|$. These inequalities follow by applying the triangle inequality for norms and noting that $\|A\|_{2}=\|A^{T}\|_{2}$ and $\|A\|=\|A^{T}\|$. In the original paper [1], $\|A\|$ is the operator norm of the matrix whose entries are absolute values of entries of $A$. Reference [2] shows that the inequality indeed holds as stated in (2) with $\|A\|$ being the operator norm of $A$ itself. An explicit value for $\kappa$ is missing or at least hard to identify in [1, 2]. Theorem 3.1.4 on page 75 in [3] provides the bounds
$\Pr(\boldsymbol{\Delta}>t)\leq\left\{\begin{array}[]{cc}\exp(-\frac{t^{2}}{8\|A\|_{2}^{2}})&0<t\leq\frac{\|A\|_{2}^{2}}{\|A\|}\\ \exp(\frac{\|A\|_{2}^{2}}{8\|A\|^{2}}-\frac{t}{4\|A\|})&t>\frac{\|A\|_{2}^{2}}{\|A\|}\end{array}\right.$. (5)
For $t>\frac{\|A\|_{2}^{2}}{\|A\|}$, we have $\frac{\|A\|_{2}^{2}}{8\|A\|^{2}}-\frac{t}{4\|A\|}<-\frac{t}{8\|A\|}$. Therefore, (5) implies HWI with $\kappa=\frac{1}{8}=0.125$. More recently, Textbook [4] gives the explicit value $\kappa=\frac{1}{8}$ at the end of the proof for Theorem B.8 on page 304. This value for $\kappa$ can also be inferred form inequality (3.28) on page 120 in [5] given by
@@
It is stated in [5] that (6) holds for arbitrary symmetric matrix $A$. However, there is a mistake in the proof. Inequality $\frac{1}{2}(-\ln(1-2x)-2x)\leq\frac{x^{2}}{1-2x}$ in the middle of page 120 does not hold for $x<0$. Nonetheless, inequality (6) certainly holds when $A$ is positive-semidefinite. Loosening the bound by writing $\|A\|_{2}^{2}+\|A\|t\leq 2\max\{\|A\|_{2}^{2},\|A\|t\}$, we arrive at HWI with $\kappa=\frac{1}{8}$. For a positive-semidefinite matrix $A$, a key result in the literature on concentration of Gaussian quadratic chaos is inequality (4.1) on page 1325 in [6] due to Laurent and Massart given by
@@
for all tail parameters $t>0$. We will refer to (7) as Laurent-Massart inequality (LMI). It is notable that LMI also appears in Example 2.12 in Textbook [7]. One can show that LMI is equivalent to [^3] See Section V.
@@
We will show in Section V that LMI implies HWI at best with $\kappa=1-\frac{\sqrt{3}}{2}\approx 0.134$.
@@ I-B Contributions and main results
The main contributions in this paper are fivefold.
@@
    We visit and slightly modify the proof given in [4] for HWI. In the course of proof, Reference [4] uses the bound
@@
    for $a,b>0$. For given $0<b<1$, we determine the smallest $a$ (tightest bound) such that (10) holds. As a result, we are able to increase $\kappa$ from $0.125$ to at least $0.145$ for arbitrary symmetric matrix $A$. More precisely, we have the following proposition.
@@
    For $0<b<1$ and integer $m\geq 1$ define the sequence of functions
    $\theta_{m}(b)=-\frac{1}{b^{m+1}}\Big(\ln(1-b)+b+\frac{b^{2}}{2}+\frac{b^{3}}{3}+\cdots+\frac{b^{m}}{m}\Big).$ (11)
    If $A$ is symmetric, then HWI holds with
@@
    where $b^{*}$ is the unique solution for $b$ in the equation $2b\theta_{1}(b)=1$.
@@
    A special case of Lemma 1 in Section III shows that the smallest value for the parameter $a$ in (10) is $\theta_{1}(b)$. This optimal value for $a$ is also reported independently in Reference [8] in Corollary 4 on Page 5. The functions $\theta_{m}(b)$ will appear again in Theorem 1 and Theorem 2 in below.
@@
    In the positive-semidefinite case, we present a sharper version of LMI. To prove (7), Reference [6] relies on the bound
    $-\ln(1-x)\leq x+\frac{x^{2}}{2(1-x)},\,\,\,\,\,0\leq x\leq 1.$ (13)
@@
    $-\ln(1-x)\leq x+\frac{x^{2}}{2(1-ax)},\,\,\,\,\,0\leq x\leq b,$ (14)
    for $a,b>0$. We show that for every $\frac{2}{3}<a\leq 1$, this inequality holds for $b$ at least as large as $\frac{3a-2}{a(2a-1)}$. As a result, we derive an improved LMI as presented in the next proposition.
@@
    Choosing $a=b=1$ and $c=\infty$ in (17) recovers LMI in (8). In Subsection II-A, we minimize the upper bound on the right side of (17) in terms of $a$ and compare the resulting optimized inequality with LMI to demonstrate the improvement. We will see that the minimizing value for $a$ is the unique root inside the interval $(\frac{2}{3},1)$ of a quintic polynomial equation [^5] See Equation (46). and it depends on the matrix $A$ and the tail parameter $t$ only through the ratio $\rho=\frac{t\|A\|}{\|A\|_{2}^{2}}$. We are able to approximate this unique root by an analytic formula [^6] See Equation (41). where the approximation error is shown to be less than $0.035$ regardless of $A$ and $t$. As a consequence of Proposition 2, we offer the next result which is a stronger version of Proposition 1 in the positive-semidefinite case.
@@
    Let $A$ be positive-semidefinite. The inequality in (17) implies HWI at best with
@@
    Moreover, LMI in (7) implies HWI at best with $\kappa=1-\frac{\sqrt{3}}{2}\approx 0.134$.
@@
    We explore beyond HWI for arbitrary symmetric matrix $A$ by considering an extension of (10), i.e., an inequality of the form
@@
    for a given integer $m\geq 1$ and $a,b>0$. We show [^7] See Lemma 1 in Section III. that for every $m\geq 1$ and $0<b<1$, the smallest $a$ (tightest upper bound) for which (23) holds is $\theta_{m}(b)$ as defined in (11). Consequently, we obtain a sequence of concentration bounds for Gaussian quadratic chaos that are written in terms of Schatten norms of AA. [^8] See Subsection I.C for the definition of Schatten $p$-norm.
    Theorem 1.
    Let $m\geq 1$ be an integer, $0<b<1$ and $\|A\|_{m+1}$ be Schatten norm of order $m+1$ for a symmetric matrix $A$. Then
@@
    Choosing $m=1$ in (24) recovers HWI in (2) with $\kappa=\kappa_{1}(b)$. Schatten norm $\|A\|_{m+1}$ should not be mistaken with the $L_{p}$-norm of $A$ with $p=m+1$ unless $m=1$ in which case $\|A\|_{2}$ is Hilbert-Schmidt norm of $A$. The upper bound $e^{\Lambda_{m}(t,b)}$ in (24) is to be minimized over $0<b<1$ and the choice of $m\geq 1$. In general, the sequence of upper bounds $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ for $m=1,2,3,\cdots$ is not monotone in $m$ for given $A$ and $t$. However, we conjecture that it undergoes a phase transition as stated next.
@@
    For a given symmetric matrix $A$, there exists a critical threshold $\tau_{c}>0$ such that if $t<\tau_{c}$, then $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ is the smallest for $m=1$ and if $t>\tau_{c}$, then $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ is the smallest when $m$ grows to infinity. [^10] More precisely, for $t>\tau_{c}$, the limit $\lim_{m\to\infty}\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ exists and it is smaller than $\inf_{b\in(0,1)}e^{\Lambda_{m}(t,b)}$ for every $m\geq 1$.
    If Conjecture 1 is true, then one must always choose $m=1$ or let $m$ grow to infinity. In Subsection II-B, we will study the bounds in (24) both through numerical examples and analytically where two corollaries to Theorem 1 are stated. The first corollary looks into the limiting value of (24) as $m$ grows large. This leads to the simple bound
@@
    which we refer to as the $m_{\infty}$-bound or $m_{\infty}$-inequality. The right hand side in (27) depends on $A$ only through the dimension $n$ and its operator norm $\|A\|$. The second corollary presents an estimate (upper bound) for the critical threshold $\tau_{c}$ mentioned in Conjecture 1 in above.
@@
    The upper bounds in Theorem 1 become tighter than the upper bound in HWI only when $t>\tau_{c}$, i.e., for all sufficiently large values of $t$. We raise the question whether one can improve over HWI for all sufficiently small values of $t$. We answer this question in the affirmative. We are able to present what we refer to as a “twin inequality” to HWI which is tighter than HWI not only for all sufficiently small $t$, but also for all sufficiently large $t$. This bound depends on all norms $\|A\|,\|A\|_{2}$ and $\|A\|_{3}$.
    Theorem 2.
@@
    The bound in (31) is named a twin to HWI in the sense that $\min\{\eta_{1}(t),\eta_{2}(t)\}=\frac{t^{2}}{\|A\|_{2}^{2}}+O(t^{3})$ as $t\to 0$ and $\min\{\eta_{1}(t),\eta_{2}(t)\}=\frac{t}{\|A\|}+O(1)$ as $t\to\infty$. These statements are verified in Subsection II-C. Hence, the inequality in (31) is a look-alike to HWI both as $t\to 0$ and $t\to\infty$. Note that the constant $\kappa^{\prime}\approx 0.187$ is slightly larger than the constant $\kappa\approx 0.145$ in HWI. This is exactly why the twin bound is sharper than HWI for both small and large values of $t$.
@@
    To further explore the $m_{\infty}$-bound, we raise and attempt to answer the following question in Subsection II-D.
    Let $A$ be an $n\times n$ positive-semidefinite matrix for which we only know the dimension $n$ and the value of its operator norm $\|A\|$ denoted by $\nu$. How much is $\sup_{A\in\mathbb{S}^{n}_{+}:\|A\|=\nu}\Pr(\boldsymbol{\Delta}>t)$ in terms of $n,\nu,t$? Here, $\mathbb{S}^{n}_{+}$ is the cone of $n\times n$ positive-semidefinite matrices.
    We present five bounds on $\Pr(\boldsymbol{\Delta}>t)$ that depend on $A$ only through $n$ and $\|A\|$. One candidate is the $m_{\infty}$-bound in (27). The second and third candidates are relaxed versions of HWI and LMI, respectively, which do not depend on $\|A\|_{2}$ anymore. The relaxed HWI is
@@
    where $F_{\chi^{2}_{n}}(\cdot)$ is the cumulative distribution function (CDF) for a $\chi^{2}$ random variable with $n$ degrees of freedom. The fifth candidate is a loosened version of the $\chi^{2}$-bound known as the large deviations bound given by
    $\Pr(\boldsymbol{\Delta}>t)<\frac{1}{\sqrt{e}}\Big(\frac{e(1+\frac{t}{\|A\|})}{n}\Big)^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}},\,\,\,\,\,\frac{t}{\|A\|}\geq n-1.$ (35)
    We emphasize that (35) is only valid for $\frac{t}{\|A\|}\geq n-1$. Derivations of the relaxed HWI, the relaxed LMI, the $\chi^{2}$-bound and the large deviations bound are presented in Subsection II-D. Note that all of the bounds in (27), (32), (33), (34) and (35) depend on $A$ and $t$ only through the dimension $n$ and the ratio $\frac{t}{\|A\|}$. Our first result is the following.
@@
    The $m_{\infty}$-bound in (27) is sharper than the relaxed HWI in (32), the relaxed LMI in (33) and the large deviations bound in (35) for every positive-semidefinite matrix $A$ and $t>0$.
@@
[…]